The problem reads as follows: y+5=-(x+2)
It asks to graph the problem. Could you please show me how to put this in this in the correct form and graph it? If I can see one done, I think I can do the rest.
Thank you.........Ann

Answer by jojo14344(1513)



Given: y+5=-(x+2) ---------> y+5=-x-2
Arranging to Slope-Intercept Form, y=mx+b:
y= -x-2-5
y= -x-7, you're Line Eqn.

Let fy=0: 0=(-x)-7 ===> x= -7, X-intercept
Let fx=0: y=0-7 ===> y= -7, Y-Intercept

Let's plot:
---->

Thank you,
Jojo